OpenStudy (anonymous):

In square MNKJ, diagonal MK measures square root 50 and in rectangle IJKL, diagonal KI measures 13. What is the length of IJ or KL?

OpenStudy (anonymous):

picture with problem

OpenStudy (anonymous):

12

OpenStudy (anonymous):

how is it 12?? Can you just show me how you got the answer??

OpenStudy (anonymous):

in the square, triangle mkj is right angle triangle. pythagoras theorem tells us that MJ^2+JK^2 = MK^2 but we know that MJ = JK, since MNJK is a square. so 2JK^2 =MK^2 we know that MK^2 = 50 so JK = 5

OpenStudy (anonymous):

again, in the rectangle IJKL, IJK is a right angle triangle. you know IK and JK, find IJ

OpenStudy (anonymous):

MK^2= square root 50.....not 50

OpenStudy (anonymous):

no. read the question. MK measures sqrt 50 so MK^2 = 50

OpenStudy (anonymous):

so IJ is 12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks :)